[Koji, Jenne, Kevin]
Jenne worked on fixing REFL11 last week (see elog 4034) and was able to measure an electrical transfer function. Today, I tried to measure an optical transfer function but REFL11 is still not responding to any optical input. I tried shining both the laser and a flashlight on the PD but could not get any DC voltage.
I also completed the characterizations of POX. I redid the optical transfer function and shot noise measurements. I also took a time series of the RF output from the PD when it was powered on with no light. This measurement shows oscillations at about 225 MHz. I also measured the spectrum with no light which also shows the oscillations at 225 MHz and smaller oscillations at ~455 MHz.
The plots can be found at http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/POX?action=show.